OpenStudy (insa):

A boy has a certain number of sweets. If he eats 16 a day, they will last him 2 days longer than if he eats 18 a day. How many sweets has he?

OpenStudy (insa):

@paki

OpenStudy (phi):

first, name the "unknowns" x = # of sweets d = # of days it takes to eat all the sweets if he eats 18 each day from those definitions, we can write the equation x= 18d (this says if you eat 18 each day for "d" days, you get the total number of sweets)

OpenStudy (phi):

If he eats 16 a day, they will last him 2 days longer this means it will take d+2 days to eat all of the sweets, and we can write 16(d+2)= x now we use both of these equations: x= 18d x= 16(d+2)

OpenStudy (phi):

from those two equations we see 18d must equal 16(d+2): 18d = 16(d+2) now solve for d. distribute the 16 on the right side: 18d= 16d + 32 now add -16d to both sides. we get 2d = 32 and d= 16 now we can find x. use either of the equations, for example x= 18d x= 18*16

OpenStudy (insa):

Thankyou :)